OpenStudy (anonymous):

Which one of the following sets of numbers represents the lengths of the sides of a right triangle? A. 5, 12, 13 B. 13, 35, 37 C. 9, 40, 42 D. 6, 9, 10 <--- @iambatman @satellite73

OpenStudy (anonymous):

does \[6^2+9^2=10^2\]?

OpenStudy (anonymous):

hint: no it doesn't so it is not a right triangle try some others

OpenStudy (anonymous):

What sat showed you was the Pythagorean theorem, I'm sure you know of it?

OpenStudy (anonymous):

Sorry typo I meant C. Though that might be wrong too. :) @iambatman @satellite73

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

C is wrong as well

OpenStudy (anonymous):

5,12,13 is a well known right triangle you can check that \[5^2+12^2=13^2\]

OpenStudy (anonymous):

\[25+144=169\]